# Uniqueness of asymptotically cylindrical steady gradient Ricci solitons

### v1

@@ Abstract.
We show that the Bryant soliton is the unique asymptotically cylindrical steady gradient Ricci soliton, in any dimension $n\geq 3$ and without any curvature assumptions. This generalizes a celebrated theorem of Brendle. We also prove that any steady gradient Ricci soliton asymptotic to a cylinder over the homogeneous lens space $\mathbb{S}^{2m+1}/\mathbb{Z}_{k}=L_{m,k}$, for $m\geq 1$ and $k\geq 3$, is a noncollapsed Appleton soliton on the complex line bundle $\mathcal{O}(-k)$ over $\mathbb{CP}^{m}$. In dimension 4, our results lead to a classification of steady gradient Ricci soliton singularity models on smooth manifolds which possess a tangent flow at infinity of the form $(SU(2)/\Gamma)\times\mathbb{R}$.
@@ 1. Introduction
A Ricci soliton is a complete $n$-dimensional Riemannian manifold $(M^{n},g)$ with a complete vector field $X$ such that
(1.1) $\operatorname{Ric}=\frac{1}{2}\mathcal{L}_{X}g+\lambda g,$
where $\mathcal{L}$ is the Lie derivative, and $\lambda\in\mathbb{R}$ determines the soliton type: shrinking ($\lambda>0$), steady ($\lambda=0$), or expanding ($\lambda<0$). If $X=\nabla f$ for a function $f:M\to\mathbb{R}$, then $(M^{n},g,f)$ is called a gradient Ricci soliton and the equation (1.1) becomes
(1.2) $\operatorname{Ric}=\mathrm{Hess}_{f}+\lambda g.$
Throughout this article, we abbreviate GRS = gradient Ricci soliton.
Ricci solitons are self-similar solutions of the Ricci flow. Shrinking and steady Ricci solitons are special instances of ancient solutions, which are solutions of Ricci flow defined for all negative time. Ancient solutions arise as blowup limits of Ricci flow at singularities; we call these blowup limits singularity models. Classifying Ricci solitons and ancient solutions provides a way to understand singularity models, which is in turn crucial for understanding singularity formation in Ricci flow.
In dimension 3, Perelman [perelman1] showed that finite-time singularity models of Ricci flow are ancient $\kappa$-solutions: non-flat, $\kappa$-noncollapsed ancient solutions with bounded and nonnegative sectional curvature. The classification of ancient $\kappa$-solutions in dimension 3 was completed by Brendle [brendle-acta], building on techniques developed in his work on 3-dimensional steady GRSs [brendle2013rotational] and the foundations laid by many other authors. For more information on Ricci solitons and ancient solutions in dimensions $\leq 3$, we refer to the books [chow2023ricci, chowluni] and references therein.
In dimensions $\geq 4$, ancient $\kappa$-solutions no longer capture all singularity models of Ricci flow. In fact, singularity models in general need not enjoy any curvature bound unless one imposes strong positivity conditions on the initial data, such as positive isotropic curvature in dimension 4 [hamilton-pic]. Yet, all known classification results for Ricci solitons and ancient solutions in dimensions $\geq 4$ impose curvature hypotheses alongside other structural assumptions; see for instance [brendle2014rotational, brendle-naff, zz, deng-zhu, chodosh, chodosh-fong, cds]. This raises the natural question: can one classify Ricci solitons and/or ancient solutions in dimensions $\geq 4$ without imposing any curvature assumptions?
In this paper, we answer this question affirmatively for steady GRSs that are asymptotic to cylinders over certain spherical space forms. The notion of asymptotic cylindricity is standard and arises naturally in various situations; see e.g. [brendle2013rotational, deng-zhu, bamler-3].
@@
Let $n\geq 3$. Let $\mathcal{S}=\mathbb{S}^{n-1}/\Gamma$ be a spherical space form of dimension $n-1$. A steady GRS $(M^{n},g,f)$ is said to be asymptotically $\mathcal{S}$-cylindrical if the following holds. Let $\Phi_{t}$ be the flow of the vector field $\nabla f$. Take an arbitrary sequence $p_{m}$ of marked points going to infinity and consider the rescaled flows
@@ 1.1. Uniqueness of the Bryant soliton
A famous theorem of Brendle [brendle2013rotational, brendle2014rotational] states that any asymptotically cylindrical steady GRS of dimension $n\geq 3$ with positive sectional curvature must be rotationally symmetric, hence isometric to the Bryant soliton [bryant]. In three dimensions, the hypotheses are met for any non-flat, $\kappa$-noncollapsed steady GRS, and Brendle’s result in this setting resolved a major conjecture left open in Perelman’s first paper [perelman1]. Ten years later, Zhao and Zhu [zz] generalized Brendle’s theorem in dimensions $\geq 4$ by relaxing the positive sectional curvature condition to the weaker condition $2P<R$, where $P$ is the operator norm of the Riemann curvature acting on symmetric 2-tensors.
Our first result proves uniqueness of the Bryant soliton in a much broader setting. We show that uniqueness holds in any dimension only assuming asymptotic cylindricity, without any curvature hypotheses.
@@
Let $(M^{n},g,f)$ be an asymptotically cylindrical steady GRS of dimension $n\geq 3$. Then $(M,g,f)$ is the Bryant soliton (i.e. they are isometric up to scale).
In dimension 4, Deng and Zhu [deng-zhu]*Theorem 1.5 established a sufficient condition for a steady GRS to be asymptotically cylindrical. Combining Theorem 1.4 with their result, we get:
@@
Any 4-dimensional $\kappa$-noncollapsed steady GRS with positive Ricci curvature and linear curvature decay must be the Bryant soliton.
This affirmatively answers a conjecture posed by Deng and Zhu in the same paper. Corollary 1.5 is comparable to Brendle’s 4-dimensional result [brendle2014rotational], which reaches the same conclusion assuming positive isotropic curvature instead of ‘positive Ricci curvature and linear curvature decay’.
@@
There is no asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRS for any $n\geq 3$.
@@ 1.2. Uniqueness of Appleton’s noncollapsed solitons
In [appleton], Appleton discovered a new family of $\kappa$-noncollapsed steady GRSs in each even dimension $\geq 4$, and indicated numerical evidence of these solitons arising as Ricci flow singularity models. A uniqueness theorem is therefore in order, and this is our next result. To state this, we first review the construction of Appleton’s solitons.
@@
    Each $(M,g,f)\in\mathcal{A}_{m,k}$ is $U(m+1)$-invariant, $\kappa$-noncollapsed, and asymptotically $L_{m,k}$-cylindrical (see Appendix A).
@@
Let $(M,g,f)$ be an asymptotically $L_{m,k}$-cylindrical steady GRS with $m\geq 1$ and $k\geq 3$. Then $(M,g,f)\in\mathcal{A}_{m,k}$.
@@
1.3. 4D steady GRSs with homogeneous 3-cylindrical tangent flows
Any 3-dimensional spherical space form that is homogeneous with the round metric (i.e. its isometry group acts transitively) is of the form $SU(2)/\Gamma$ for some finite subgroup $\Gamma\subset SU(2)$ acting from the right. Every finite subgroup of $SU(2)$ is one of the following up to conjugacy (see e.g. [cmp-su2]):
@@
We have $L_{1,k}\cong SU(2)/\mathbb{Z}_{k}$. Thus, Theorems 1.4, 1.6, and 1.8 together classify the asymptotically $SU(2)/\mathbb{Z}_{k}$-cylindrical steady GRSs for $k\geq 1$. To complete the picture, we classify the asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs for all other $\Gamma$. For $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$, no such solitons exist:
@@
There is no asymptotically $SU(2)/\Gamma$-cylindrical steady GRS for $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$.
For $\Gamma=\tilde{D}_{k}$ ($k\geq 2$), the story is different. Let $E_{k}$ be the non-orientable 2-plane bundle over $\mathbb{RP}^{2}$ with Euler number ±$\pm k$. The sphere bundles of $E_{k}$ are diffeomorphic to $SU(2)/\tilde{D}_{k}$, and there is a double covering $\mathcal{O}_{\mathbb{CP}^{1}}(-2k)\to E_{k}$. We will see that any Appleton soliton $A\in\mathcal{A}_{1,2k}$ descends through the map $\mathcal{O}_{\mathbb{CP}^{1}}(-2k)\to E_{k}$, giving an asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRS on $E_{k}$ denoted by $A/\mathbb{Z}_{2}$. Moreover, we show that $A/\mathbb{Z}_{2}$ are the only steady GRSs with such asymptotics.
@@
Let $(M,g,f)$ be an asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRS with $k\geq 2$. Then $(M,g,f)=A/\mathbb{Z}_{2}$ for some $A\in\mathcal{A}_{1,2k}$.
Combining the 4-dimensional versions of our results (Theorems 1.4, 1.6, 1.8, 1.10, and 1.11), we arrive at the following classification.
Corollary 1.12.
Let $(M^{4},g,f)$ be an asymptotically $\mathcal{S}$-cylindrical steady GRS, where $\mathcal{S}=SU(2)/\Gamma$ is a homogeneous 3-dimensional spherical space form with $\Gamma\subset SU(2)$ finite. Then exactly one of the following holds:
  • $\Gamma=\{1\}$, and $(M,g,f)$ is the Bryant soliton.
  • $\Gamma=\mathbb{Z}_{k}$ for some $k\geq 3$, and $(M,g,f)\in\mathcal{A}_{1,k}$.
  • $\Gamma=\tilde{D}_{k}$ for some $k\geq 2$, and $(M,g,f)=A/\mathbb{Z}_{2}$ for some $A\in\mathcal{A}_{1,2k}$.
We highlight one further consequence of this. Bamler et al. [bamler-3] showed that any tangent flow at infinity of a 4-dimensional steady GRS singularity model is one of
(1.12) $\mathbb{R}^{4}/\Gamma\text{ (but not }\mathbb{R}^{4}\text{)},\quad(\mathbb{S}^{3}/\Gamma)\times\mathbb{R},\quad\mathbb{S}^{2}\times\mathbb{R}^{2},\quad((\mathbb{S}^{2}\times\mathbb{R})/\mathbb{Z}_{2})\times\mathbb{R},$
where $\Gamma$ is a finite subgroup of $O(4)$. Moreover, if one tangent flow at infinity is $(\mathbb{S}^{3}/\Gamma)\times\mathbb{R}$, then the soliton is asymptotically $\mathbb{S}^{3}/\Gamma$-cylindrical. Hence, Corollary 1.12 classifies all steady GRS singularity models with a tangent flow at infinity of the form $(SU(2)/\Gamma)\times\mathbb{R}$, where $\Gamma$ is a finite subgroup of $SU(2)$.
@@ 1.4. Summary of results
For the reader’s convenience, we gather our classification results into Table 1 below.
@@
$SU(2)/\Gamma$ ($\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$) None 1.10
$SU(2)/\tilde{D}_{k}$ ($k\geq 2$) $A/\mathbb{Z}_{2}$ for $A\in\mathcal{A}_{1,2k}$ 1.11
Table 1. Classification of asymptotically $\mathcal{S}$-cylindrical steady GRSs for various spherical space forms $\mathcal{S}$.
We also have Corollary 1.5 and Corollary 1.12 as consequences in dimension 4.
1.5. Remarks on the orbifold case
While this paper exclusively works with smooth Riemannian manifolds, a robust singularity theory of Ricci flow in dimensions $\geq 4$ requires objects of lower regularity. Indeed, Bamler’s compactness theory [bamler-a, bamler-b, bamler-c] shows that finite-time singularity models of Ricci flow are orbifold Ricci solitons with singular locus of codimension $\geq 4$.
It is therefore of interest to see if our results generalize to the orbifold setting, and we believe they do. Our proofs use local properties like the maximum principle and the analyticity of Ricci solitons, which generalize directly to orbifolds as one can pass to local uniformizing covers and apply results for smooth manifolds there. We also use the standard formalism for cohomogeneity one manifolds, of which a version for orbifolds exists [gonzalez-alvaro-phd, borzellino-phd]. To this end, the only major change occurs in Proposition 4.2(d), where $K/H$ is now allowed to be a spherical space form $\mathbb{S}^{\ell}/\Gamma$. Naively applying this change to our proofs, it appears that quotients of the Bryant soliton are the only new possible steady GRSs that arise in the expanded category of smooth Riemannian orbifolds. Hence, we conjecture:
Conjecture 1.13.
Table 1 correctly classifies asymptotically $\mathcal{S}$-cylindrical steady GRSs on smooth Riemannian orbifolds for certain $\mathcal{S}$, provided that one includes the appropriate Bryant soliton quotient as an extra possibility in each row.
Verifying this conjecture requires checking if global and asymptotic properties of steady GRSs on Riemannian manifolds carry over to orbifolds. For instance, we used connectedness at infinity (Theorem 2.4) whose proof does not generalize directly to orbifolds. Instead, this property is separately established in [bamler-3]*Appendix A for 4-dimensional steady GRS singularity models with a tangent flow at infinity of the form $(\mathbb{S}^{3}/\Gamma)\times\mathbb{R}$.
In a recent article [deng-orbifold], Deng proved that any asymptotically $\mathcal{S}$-cylindrical steady GRS on a smooth orbifold with positive sectional curvature must be a quotient of the Bryant soliton. Unlike for manifolds, it is impossible to drop the curvature hypothesis (at least entirely), as the solitons in $\mathcal{A}_{m,k}$ would furnish counterexamples. The reason our methods cannot apply here is because Deng employs a reduction to manifolds that fails without curvature conditions. Specifically, Deng obtains a covering of the orbifold GRS $M$ by an asymptotically cylindrical steady GRS on a manifold $\tilde{M}$ with positive sectional curvature. Then Brendle’s uniqueness result [brendle2014rotational] (alternatively Theorem 1.4) applies to $\tilde{M}$. However, the existence of such a cover requires the singular locus of $M$ to consist of a single point, which demands at least the positivity of Ricci curvature (see [deng-orbifold]*§4).

### v2

@@ Abstract.
We show that the Bryant soliton is the unique asymptotically cylindrical steady gradient Ricci soliton, in any dimension $n\geq 3$ and without any curvature assumptions. This generalizes a celebrated theorem of Brendle. We also prove that any steady gradient Ricci soliton asymptotic to a cylinder over the homogeneous lens space $\mathbb{S}^{2m+1}/\mathbb{Z}_{k}=L_{m,k}$, for $m\geq 1$ and $k\geq 3$, is a noncollapsed Appleton soliton on the complex line bundle $\mathcal{O}(-k)$ over $\mathbb{CP}^{m}$. Specializing to dimension 4, we classify steady gradient Ricci soliton singularity models on smooth orbifolds with tangent flows at infinity of the form $(SU(2)/\Gamma)\times\mathbb{R}$.
@@ 1. Introduction
Let $(M^{n},g)$ be a complete $n$-dimensional smooth Riemannian manifold or orbifold, and let $f:M\to\mathbb{R}$ be a smooth function. The triple $(M^{n},g,f)$ is called a gradient Ricci soliton (GRS) if
(1.1) $\operatorname{Ric}=\mathrm{Hess}_{f}+\lambda g,$
where $\lambda\in\mathbb{R}$ determines the soliton type: shrinking ($\lambda>0$), steady ($\lambda=0$), or expanding ($\lambda<0$).
GRSs generate self-similar solutions of the Ricci flow. Shrinking and steady GRSs are special instances of ancient solutions, which are solutions of Ricci flow defined for all negative time. Ancient solutions are important in that they arise as blowup limits of Ricci flow at singularities. The classification of GRSs and ancient solutions is therefore of significant interest, and has crucial implications for Ricci flow.
In dimension 3, Perelman [perelman1] showed that finite-time singularities of Ricci flow are modeled on ancient $\kappa$-solutions: non-flat, $\kappa$-noncollapsed ancient solutions on smooth manifolds with bounded and nonnegative sectional curvature. The classification of ancient $\kappa$-solutions in dimension 3 was completed by Brendle [brendle-acta] in the noncompact case and by Brendle–Daskalopoulos–Sesum in the compact case [bds], building on Brendle’s work on 3-dimensional steady GRSs [brendle2013rotational] and the foundations laid by many other authors. For more information on GRSs and ancient solutions in dimensions $\leq 3$, we refer to the books [chow2023ricci, chowluni] and references therein.
In dimensions $\geq 4$, ancient $\kappa$-solutions no longer model all singularities of Ricci flow. In fact, blowup limits in general need not enjoy any curvature bound unless one imposes strong positivity conditions on the initial data, such as positive isotropic curvature in dimension 4 [hamilton-pic]. To complicate matters further, Bamler’s compactness and partial regularity theory [bamler-a, bamler-b, bamler-c] shows that singularity models of Ricci flow in dimensions $\geq 4$ are generally not smooth manifolds. Yet, virtually all known classification results for GRSs and ancient solutions in dimensions $\geq 4$ take place on smooth manifolds with curvature hypotheses, alongside other structural assumptions; see for instance [brendle2014rotational, brendle-naff, zz, deng-zhu, chodosh, chodosh-fong, cds].
In this paper, we classify steady GRSs that are asymptotic to cylinders over certain spherical space forms. Our results require no curvature hypotheses. In dimension 4, our results also hold for steady GRS singularity models on smooth orbifolds with isolated singularities. This is the expected regularity for singularity models of 4-dimensional Ricci flow. We use a standard notion of asymptotic cylindricity which arises naturally in various situations (see e.g. [brendle2013rotational, deng-zhu, bamler-3]):
@@
Let $n\geq 3$. Let $\mathcal{S}=\mathbb{S}^{n-1}/\Gamma$ be a spherical space form of dimension $n-1$. A steady GRS $(M^{n},g,f)$ on a smooth manifold or orbifold is called asymptotically $\mathcal{S}$-cylindrical if the following holds. Let $\Phi_{t}$ be the flow of $\nabla f$. Take an arbitrary sequence $p_{m}$ of points going to infinity and consider the rescaled flows
@@ 1.1. Uniqueness of the Bryant soliton
A famous theorem of Brendle [brendle2013rotational, brendle2014rotational] states that any asymptotically cylindrical steady GRS of dimension $n\geq 3$ on a smooth manifold with positive sectional curvature must be rotationally symmetric, hence isometric to the Bryant soliton [bryant]. In three dimensions, the hypotheses are met for any non-flat, $\kappa$-noncollapsed steady GRS, and Brendle’s result in this setting resolved a major conjecture left open in Perelman’s first paper [perelman1]. Ten years later, Zhao and Zhu [zz] generalized Brendle’s theorem in dimensions $\geq 4$ to hold under the weaker curvature condition $2P<R$, where $P$ is the operator norm of the Riemann curvature acting on symmetric 2-tensors. Deng [deng-orbifold] has also recently extended Brendle’s theorem to smooth orbifolds with positive sectional curvature.
We show that the curvature hypothesis in Brendle’s theorem can be entirely dropped. Thus, uniqueness of the Bryant soliton holds only assuming asymptotic cylindricity for smooth manifolds in any dimension.
@@
Let $(M^{n},g,f)$ be an asymptotically cylindrical steady GRS on a smooth manifold of dimension $n\geq 3$. Then $(M,g,f)$ is the Bryant soliton (i.e. they are isometric up to scale).
In dimension 4, Deng and Zhu [deng-zhu]*Theorem 1.5 established conditions for a steady GRS on a smooth manifold to be asymptotically cylindrical. They conjectured that such a soliton must be the Bryant soliton. Theorem 1.4 confirms this:
@@
Any 4-dimensional $\kappa$-noncollapsed steady GRS on a smooth manifold with positive Ricci curvature and linear curvature decay must be the Bryant soliton.
Corollary 1.5 is comparable to Brendle’s 4-dimensional result [brendle2014rotational], which reaches the same conclusion under the assumption of positive isotropic curvature.
@@
There is no asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRS on a smooth manifold for any $n\geq 3$.
@@ 1.2. Uniqueness of Appleton’s noncollapsed solitons
In [appleton], Appleton discovered a new family of $\kappa$-noncollapsed steady GRSs in each even dimension $\geq 4$, and indicated numerical evidence that these solitons do arise as Ricci flow blowup limits. A uniqueness theorem is therefore in order, and this is our next result. To state this, we first review the construction of Appleton’s solitons.
@@
    Each $(M,g,f)\in\mathcal{A}_{m,k}$ is $U(m+1)$-invariant, $\kappa$-noncollapsed, and asymptotically $L_{m,k}$-cylindrical.
@@
Let $(M,g,f)$ be an asymptotically $L_{m,k}$-cylindrical steady GRS on a smooth manifold, with $m\geq 1$ and $k\geq 3$. Then $(M,g,f)\in\mathcal{A}_{m,k}$.
@@
1.3. Asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs
Any 3-dimensional spherical space form that is homogeneous with the round metric (i.e. its isometry group acts transitively) must be $SU(2)/\Gamma$ for some finite subgroup $\Gamma$ acting from the right. The finite subgroups of $SU(2)$ are classified up to conjugacy (see e.g. [cmp-su2]):
@@ 1.3. Asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs
We have $L_{1,k}\cong SU(2)/\mathbb{Z}_{k}$. Thus, Theorems 1.4, 1.6, and 1.8 classify the asymptotically $SU(2)/\mathbb{Z}_{k}$-cylindrical steady GRSs for $k\geq 1$. Using the results of Bamler et al. [bamler-3], these theorems generalize to 4-dimensional steady GRSs on smooth orbifolds with isolated singularities which are singularity models in the sense of [bamler-3]*Definition 1.1. One simply adds in the appropriate Bryant soliton quotients:
Corollary 1.10.
Let $(M^{4},g,f)$ be a steady GRS singularity model on a smooth orbifold with isolated singularities. Assume $(M,g,f)$ is asymptotically $SU(2)/\mathbb{Z}_{k}$-cylindrical.
  1. (a)
    If $k=1$, then $(M,g,f)$ is the Bryant soliton.
  2. (b)
    If $k=2$, then $(M,g,f)$ is the Bryant soliton quotiented by $\mathbb{Z}_{2}$.
  3. (c)
    If $k\geq 3$, then either $(M,g,f)\in\mathcal{A}_{m,k}$, or $(M,g,f)$ is the Bryant soliton quotiented by $\mathbb{Z}_{k}$.
Meanwhile, for $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$, we will prove the following uniqueness result:
@@
Let $(M^{4},g,f)$ be a steady GRS singularity model on a smooth orbifold with isolated singularities. If $(M,g,f)$ is asymptotically $SU(2)/\Gamma$-cylindrical, where $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$, then $(M,g,f)$ is the Bryant soliton quotiented by $\Gamma$.
For the remaining case $\Gamma=\tilde{D}_{k}$ ($k\geq 2$), the story is different. Each $A\in\mathcal{A}_{1,2k}$ has a $\mathbb{Z}_{2}$-quotient $A/\mathbb{Z}_{2}$, which is asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical and lives on the non-orientable 2-plane bundle over $\mathbb{RP}^{2}$ with Euler number ±$\pm k$. See §7.3 for details. We prove that $A/\mathbb{Z}_{2}$ are the only steady GRSs with such asymptotics aside from Bryant soliton quotients.
@@
Let $(M^{4},g,f)$ be a steady GRS singularity model on a smooth orbifold with isolated singularities. If $(M,g,f)$ is asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical for some $k\geq 2$, then $(M,g,f)$ is either $A/\mathbb{Z}_{2}$ for some $A\in\mathcal{A}_{1,2k}$, or it is the Bryant soliton quotiented by $\tilde{D}_{k}$.
@@ 1.4. Summary of results
Table 1 summarizes our classification results for asymptotically $\mathcal{S}$-cylindrical steady GRSs on smooth manifolds.
@@
$SU(2)/\Gamma$ ($\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$) None 1.11
$SU(2)/\tilde{D}_{k}$ ($k\geq 2$) $A/\mathbb{Z}_{2}$ for $A\in\mathcal{A}_{1,2k}$ 1.12
Table 1. Classification of asymptotically $\mathcal{S}$-cylindrical steady GRSs on smooth manifolds for various spherical space forms $\mathcal{S}$.
$\Gamma\subset SU(2)$ Asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs Theorem
$\mathbb{Z}_{1}=\{1\}$ Bryant soliton 1.10
$\mathbb{Z}_{2}$ Bryant$/\mathbb{Z}_{2}$ 1.10
$\mathbb{Z}_{k}$ ($k\geq 3$) Bryant$/\mathbb{Z}_{k}$, and Appleton solitons in $\mathcal{A}_{1,k}$ 1.10
$\tilde{T}$, $\tilde{O}$, or $\tilde{I}$ Bryant$/\Gamma$ 1.11
$\tilde{D}_{k}$ ($k\geq 2$) Bryant$/\tilde{D}_{k}$, and $A/\mathbb{Z}_{2}$ for $A\in\mathcal{A}_{1,2k}$ 1.12
Table 2. Classification of asymptotically $SU(2)/\Gamma$-cylindrical steady GRS singularity models on smooth orbifolds with isolated singularities.
Table 2 classifies asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs which are singularity models on smooth orbifolds with isolated singularities. It also completely classifies such steady GRSs with a tangent flow at infinity of the form $(SU(2)/\Gamma)\times\mathbb{R}$, since this implies asymptotic $SU(2)/\Gamma$-cylindricity by [bamler-3].